OpenStudy (anonymous):

Find the x-coordinates of the points of intersection of the curves y = x^2 + 1 and y = 2x^2 - 3.

OpenStudy (amistre64):

this is where those system of equations work from algebra should kick in :)

OpenStudy (angela210793):

intersection : x^2+1=2x^2-3 x^2+1-2x^2+3=0 -x^2=-4 x^2=4 x=sqr4 x=+-2

OpenStudy (amistre64):

y = x^2 + 1 ; *-1 y = 2x^2 - 3 -y =-x^2 -1 y = 2x^2 - 3 ------------ 0 = x^2 -4 ; to wit, x^2 has to be 4. +- 2 would be appropriate